# Diffusion Posterior Proximal Sampling for Image Restoration

### v1

@@ Appendix C Additional Ablation Studies
27.29 0.182 23.11 25.65 0.240 29.04
27.58 0.179 22.81 25.15 0.249 28.32
27.03 0.169 19.01 24.65 0.221 25.92
27.47 0.168 18.98 24.97 0.217 24.90
C.2 Compute Resource.
To examine the impact of the variable $n$ on computational efficiency, we conducted an evaluation that consisted of measuring memory consumption and computational time required for the SR×$\times 4$ task. This evaluation was performed on a single NVIDIA RTX 3090 GPU, and the results are presented in Table 5, accompanied by a graphical representation in Figure 6. Notably, the computational time exhibits negligible growth with an increase in $n$, while concurrently manifesting a notable enhancement in performance. The memory consumption only changes a bit when $n=100$, but this configuration is not recommended, as it yields only modest performance gains. Our recommended setting, $n=20$, substantially enhances the image quality (13.6%) with almost negligible additional computational resources, demonstrating the efficacy of our method in achieving a harmonious trade-off between performance and computational overhead.
Table 5: Comparision of computational resources and performance with different $n$ on the SR×$\times 4$ task.
Dataset Memory / Increased (%) Speed / Increased (%) LPIPS / Gain (%)
$n$=1 2841 MB /  0.0 % 56.36 s /  0.0 % 0.235 /  0.0 %
$n$=2 2841 MB /  0.0 % 56.64 s /  0.5 % 0.221 /  6.0 %
$n$=10 2841 MB / 0.0 % 56.82 s /  0.8 % 0.207 /  11.9 %
$n$=20 2841 MB /  0.0 % 57.25 s /  1.5 % 0.203 /  13.6 %
$n$=100 3591 MB / 26.4 % 61.91 s /  9.8 % 0.201 /  14.5 %
C.3 Less Error Accumulation
As discussed in Section 5.3, sampling in proximity to the measurement yields less error accumulation. This is achieved by treating the injected noise as an adaptive correction. To verify this, LABEL:fig:priordis reports the Frobenius norm between true values $\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}$ and the predicted values $\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$.
@@ D.1 Comparison Methods
D.2 Metrics
The metrics employed for the comparison encompass: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) as distortion metrics; Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) distance, and Frechet Inception Distance (FID) (Heusel et al., 2017) as perceptual metrics.
D.3 Parameter Setting
@@
Table 6: Hyper-parameter $\lambda_{t}$ for each problem setting.
@@
D.4 Reproducibility
We provide the detailed algorithm that can be easily replicated in Algorithm 2, and the source code will be made public upon publication.
@@ Appendix E Limitations and Future Work
Our approach notably enhances perceptual metrics, yet it demonstrates a less substantial improvement in distortion metrics and, in certain tasks, experiences a slight decline. While this observation aligns with the perception-distortion trade-off phenomena as described in the literature (Blau & Michaeli, 2018), we acknowledge it as a noteworthy issue that warrants further investigation in our subsequent studies. In addition, subsequent work of this paper aims to extend the application to diverse domains, including but not limited to medical image reconstruction.
@@ Appendix F More Visual Results
In this section, we provide supplementary visual results to show the effectiveness of our proposed method. Figures 7 and 8 indicate that our method produces images with better details and quality as $n$ increases. Figure 9 to Figure 12 show the robustness of our method across different random seeds, in line with the claims made in the paper.
Figure 7: Qualitative results to illustrate the effectiveness of our proposed method and the impact of $n$ on SR ($\times$4) task with $\sigma_{y}=0.01$.
Figure 8: Qualitative results to illustrate the effectiveness of our proposed method and the impact of $n$ on SR ($\times$4) task with $\sigma_{y}=0.01$.
Figure 9: Qualitative inpainting results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
Figure 10: Qualitative SR ($\times$4) results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
Figure 11: Qualitative motion deblurring results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
Figure 12: Qualitative Gaussian deblurring results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.

### v2

@@ Appendix C Additional Experimental Results
More visual comparison of inpainting and deblurring on ImageNet.
C.1. Less Error Accumulation
As discussed in Section 5.3, sampling in proximity to the measurement yields less error accumulation. This is achieved by treating the injected noise as an adaptive correction. To verify this, Figure 9 reports the Frobenius norm between true values $\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}$ and the predicted values $\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$.
@@ C.1. Less Error Accumulation
Figure 9. We report the average value of $\|\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}-\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})\|_{F}$ on SR ($\times$4) task. The results show that our method achieves better predictive accuracy and reduces error accumulation.
[…]
We report the average value of $\|\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}-\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})\|_{F}$ on SR ($\times$4) task. The results show that our method achieves better predictive accuracy and reduces error accumulation.
@@ D.1. Comparison Methods
LGD-MC:  We follow the implementation of the algorithm in (Song et al., 2023c) to use a Monte Carlo estimate of the gradient correction to amend the denoising process. The number of Monte Carlo samples is set to 20.
@@
D.2. Parameter Setting
@@
Table 5. Hyper-parameter $\lambda_{t}$ for each problem setting.
@@
D.3. Parameter Size and Speed
The results for computational time and parameter sizes on the FFHQ model are presented in Table 6. The parameter sizes of the diffusion-based methods remain consistent as the same model was used.
Table 6. Computational time and parameter size comparison
Method PnP-ADMM Score-SDE MCG DDRM DPS DiffPIR Ours
Parameter (M) 0.56 93.56 93.56 93.56 93.56 93.56 93.56
Speed (s) 1.71 28.78 56.41 4.79 56.36 3.40 57.25
D.4. Source Code
Our implementation is now available at https://github.com/74587887/DPPS_code.
@@ Appendix E Limitations and Future Work
Our approach notably enhances perceptual metrics, yet it demonstrates a less substantial improvement in distortion metrics and, in certain tasks, experiences a slight decline. While this observation aligns with the perception-distortion trade-off phenomena as described in the literature (Blau and Michaeli, 2018), we acknowledge it as a noteworthy issue that warrants further investigation in our subsequent studies. In addition, subsequent work of this paper aims to extend the application to diverse domains, including but not limited to medical image reconstruction.
@@ Appendix F More Visual Results
In this section, we provide supplementary visual results to show the effectiveness of our proposed method. Figures 10 and 11 indicate that our method produces images with better details and quality as $n$ increases. Figure 12 to Figure 15 show the robustness of our method across different random seeds, in line with the claims made in the paper. Figure 8 provides one more visual comparison for inpainting and Gaussian deblurring tasks.
Figure 10. Qualitative results to illustrate the effectiveness of our proposed method and the impact of $n$ on SR ($\times$4) task with $\sigma_{y}=0.01$.
[…]
Figure 11. Qualitative results to illustrate the effectiveness of our proposed method and the impact of $n$ on SR ($\times$4) task with $\sigma_{y}=0.01$.
[…]
Figure 12. Qualitative inpainting results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
[…]
Figure 13. Qualitative SR ($\times$4) results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
[…]
Figure 14. Qualitative motion deblurring results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
[…]
Figure 15. Qualitative Gaussian deblurring results (Left FFHQ, Right ImageNet) with $\sigma_{y}=0.01$.
[…]